Aslam's age is half of his father's age but 15 years ago his age was just 1/3rd of father's age. find his present age now.
Answer by ewatrrr(24785)

x father's age
x - 15 = 3(x/2 - 15)
x - 15 = (3/2)x - 45
30= x/2
60 = x, father's age now. Son is 30 now
And...checking
15 years ago..
45 = 3(15)